Rare B Decays from CLEO

Karen Lingel SLAC

15 May 97

Frontiers in Contemporary Physics, Vanderbilt University, Nashville

Outline

• Introduction to CLEO • B → Kπ and ππ

• B → ρh, K^{∗}h, and ϕh • B → ωh

• B → ηh

• B → η′h

• B → η′X_{s} inclusive • Summary

CLEO II Experiment

Dataset:

• 3.3 M BB̄

• 4.8 M cc̄

• 3.9 M τ^{+}τ^{−}

Tricks of the Trade

• Beam energy constraint: ⟹ ΔE ≡ E_{B} − E_{beam} ≈ 0

• Beam-constrained mass: ⟹ Use E_{beam} to calculate mass,

M_{B} = \sqrt{E_{beam}^{2} − p^{2}}

• Event Shape ⟹ BB̄ events are “spherical” light quark (qq̄) are “jetty”.

The Virtual Calorimeter

• Characterization: the distribution of particles around the candidate axis.

10^{∘} conic segments

Sum all momenta (charged and neutral) in segment

Distinguishing features:

• 9 momentum sums

• B candidate flight direction

• Candidate axis direction Form a Fisher Discriminant with the 11 variables:

F = ∑_{i=1}^{11} α_{i}X_{i}

F 6

Particle ID

• Use dE/dx for particle ID

• Low momentum: easy!

• High momentum (p > 2 GeV): K − π separation of ≈ 1.7σ measured using tagged, high momentum daughters of D^{∗+} decays

Rare B Decays

8

B → π^{+}π^{−}, K^{+}π^{−}

• Tree Amplitudes  • Penguin Amplitudes
π^{+}π^{−} ∝ V_{ub}  π^{+}π^{−} ∝ V_{td}
K^{+}π^{−} ∝ V_{ub} sin θ_{Cabibbo}  K^{+}π^{−} ∝ V_{ts}

Previous result (PRD 53, 1039 (1996)):

B(B^{0} → π^{+}π^{−} + K^{+}π^{−}) = (1.8_{−0.5−0.3}^{+0.6+0.2} ± 0.2) × 10^{−5} What’s new:

• 30% more data

• Reoptimized analysis ⇒ 20% higher efficiency

• Tinkered with event shape variables

• Maximum likelihood fit to: M_{B}, ΔE, F, dE/dx_{1}, dE/dx_{2}, cos θ_{B}

N_{Kπ} = 21.7_{−6.0}^{+6.8} (5.6σ) N_{ππ} = 10.0_{−5.1}^{+6.0} (2.2σ)

Preliminary:

\begin{aligned}B(B^{0} → K^{+}π^{−}) & = (1.5_{−0.4−0.1}^{+0.5+0.1} ± 0.1) × 10^{−5} \\ B(B^{0} → π^{+}π^{−}) & < 1.5 × 10^{−5} \\ B(B^{0} → K^{+}K^{−}) & < 0.43 × 10^{−5} \end{aligned}

B^{+} → π^{+}π^{0}, K^{+}π^{0}

Preliminary:

\begin{aligned}N(π^{+}π^{0}) = 11.3_{−5.2}^{+6.3} & (2.8σ) & B(B^{+} → π^{+}π^{0}) & < 2.0 × 10^{−5} \\ N(K^{+}π^{0}) = 8.7_{−4.2}^{+5.3} & (2.7σ) & B(B^{+} → K^{+}π^{0}) & < 1.6 × 10^{−5} \\ N(h^{+}π^{0}) = 20.0_{−5.9}^{+6.8} & (5.5σ) & B(B^{+} → h^{+}π^{0}) & = (1.6_{−0.5−0.2}^{+0.5+0.2} ± 0.1) × 10^{−5} \end{aligned}

B^{+} → π^{+}K^{0}, K^{+}K^{0}

Preliminary:

\begin{aligned}N(π^{+}K_{S}^{0}) = 9.2_{−3.8}^{+4.3} & (3.2σ) & B(B^{+} → π^{+}K^{0}) & = (2.3_{−1.0−0.2}^{+1.1+0.2} ± 0.2) × 10^{−5} \\ N(K^{+}K_{S}^{0}) = 0.5_{−0.5}^{+3.8} & (0.2σ) & B(B^{+} → K^{+}K^{0}) & < 1.6 × 10^{−5} \\ N(h^{+}K_{S}^{0}) = 9.8_{−4.0}^{+4.5} & (4.4σ) & B(B^{+} → h^{+}K^{0}) & = (2.4_{−1.0−0.2}^{+1.1+0.2} ± 0.2) × 10^{−5} \end{aligned}

B^{0} → K^{0}K̄^{0}

Preliminary:

N_{obs} = 0 N_{bkg} = 0.6 ± 0.2

B(B^{0} → K^{0}K̄^{0}) < 1.7 × 10^{−5}

ππ, Kπ, and CP violation

• Direct CP violation (interference between two amplitudes):

⋆ Straightforward in self-tagging modes like Kπ ⋆ but extraction of CKM phases is clouded by strong phases

• Indirect CP violation (interference through B^{0} mixing):

⋆ π^{+}π^{−} is the golden mode ... ⋆ ... but interference from penguin diagram gives “penguin pollution” ⋆ ⇒ use π^{±}π^{0}, π^{0}π^{0}, and isospin relations to get around penguin

pollution.

• Amplitude triangles (relating various branching fractions) proposed as additional method for extracting CKM phases ...

• ⇒ These become amplitude quadrangles if electroweak penguins are non-negligible.

Modes with ρ, K^{∗}, ϕ, ω, η, η′

Extra handles:

• resonance mass ⋆ ρ → ππ ⋆ K^{∗} → K^{+}π, K_{S}^{0}π ⋆ ϕ → K^{+}K^{−} ⋆ ω → π^{+}π^{−}π^{0} ⋆ η → π^{+}π^{−}π^{0}, γγ ⋆ η′ → ηπ^{+}π^{−}; η → γγ

• in P → V P: helicity angle

• in ω → π^{+}π^{−}π^{0}: decay amplitude • clean modes: loosen event shape cuts Rare Hadronic Modes with ρ, K^{∗}, ϕ

17

B^{+} → ωh^{+}

Preliminary:

\begin{matrix}N(ωπ^{+}) = 9.5_{−4.2}^{+5.3} & (3.3σ) & B(B^{+} → ωπ^{+}) = (1.2_{−0.5}^{+0.7} ± 0.2) × 10^{−5} \\ N(ωK^{+}) = 8.6_{−3.9}^{+4.9} & (2.9σ) & B(B^{+} → ωK^{+}) = (1.2_{−0.5}^{+0.7} ± 0.2) × 10^{−5} \end{matrix}

B^{0} → ωρ^{0}  B^{0} → ωK^{∗0}  B^{+} → ωK^{∗+}

Preliminary:

Mode  B^{0} → ωρ^{0}  B^{0} → ωK^{∗0}  B^{+} → ωK^{∗+}
N_{obs}  0  1  0
N_{bkg}  1.2 ± 0.3  1.7 ± 0.4  0.9 ± 0.3
UL (90% CL)  < 3.4 × 10^{−5}  < 3.8 × 10^{−5}  < 11 × 10^{−5}

B^{+} → ηh^{+}

η → π^{+}π^{−}π^{0} η → γγ

Preliminary:

η → π^{+}π^{−}π^{0}  η → γγ
N_{obs}  0  0
N_{bkg}  0.7 ± 0.2  1.5 ± 0.4

B(B^{+} → ηh^{+}) < 0.8 × 10^{−5} (90% CL)

B^{0} → ηρ^{0}  B^{0} → ηK^{∗0}  B^{+} → ηK^{∗+}

Preliminary:

Mode  B^{0} → ηρ^{0}  B^{0} → ηK^{∗0}  B^{+} → ηK^{∗+}
submode η → π^{+}π^{−}π^{0} η → π^{+}π^{−}π^{0}  η → γγ  η → π^{+}π^{−}π^{0}
N_{obs}  1  0  1  0
N_{bkg}  0.8 ± 0.2  0.5 ± 0.2  0.8 ± 0.2  0.5 ± 0.2
UL (90% CL) < 8.4 × 10^{−5}  < 3.3 × 10^{−5}  < 24 × 10^{−5}

B^{+} → η′h^{+}

Preliminary:

\begin{aligned}N(η′K^{+}) = 12_{−3.4}^{+4.1} & (5.5σ) & B(B^{+} → η′K^{+}) & = (7.8_{−2.2}^{+2.7} ± 1.0) × 10^{−5} \\ N(η′π^{+}) = 1.4_{−1.1}^{+2.0} & (2.0σ) & B(B^{+} → η′π^{+}) & < 4.4 × 10^{−5} \end{aligned}

B^{0} → η′K^{∗0} B^{+} → η′K^{∗+}

Preliminary:

Mode  B^{0} → η′K^{∗0}  B^{+} → η′K^{∗+}
N_{obs}  0  0
N_{bkg}  0.32 ± 0.11  0.32 ± 0.11
UL (90% CL)  < 9.9 × 10^{−5}  < 29 × 10^{−5}

B → η′X_{s} Inclusive (in Progress)

• Look for high momentum η′: 2.0 < p_{η′} < 2.7 GeV; above momentum expected for most η′ from b → c decays (but we can’t yet rule out all b → c contributions.)

• Reduce background using “B reconstruction” technique:

⋆ η′ + K^{+} + n(π); n = 0–4, with at most one π^{0}

M(ηπ^{+}π^{−}) M(K^{+}n(π))

Preliminary: B(B → η′X_{s}) = (7.5 ± 1.5 ± 1.1) × 10^{−4} (2.0 < p_{η′} < 2.7 GeV).

Why is B(B^{+} → η′K^{+}) so large?

• Large η′K/ηK ratio is not surprising ⇒ interference of two penguin diagrams: (Lipkin, PLB 254 (1991))

• Possible other processes enhance rate?

⋆ “hairpin” diagram (OZI suppressed)

⋆ b → cc̄s ⟹ coupling to cc̄ component of η′

(Berkelman) ⟹ rescattering from η_{c}K^{+} (FSI) ⟹ cc̄ → g → η′ (Dunietz, et al.)

⋆ Anomalous ggη′ coupling (Atwood and Soni, hep-ph/9704357)

Prediction: B(B → η′X) = (8.2 ± 2.5) × 10^{−4} (2.0 < p_{η′} < 2.7 GeV)

27

Summary

28